OpenStudy (ashy98):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

ok so what is the first two terms of the sum? ? + ? + ?

OpenStudy (ashy98):

i have no idea

OpenStudy (anonymous):

the first two terms are f(1) + f(2) so what are they?

OpenStudy (anonymous):

ok heres a hint because there is a lot more to the problem

OpenStudy (anonymous):

the first two terms are 2 + 4

OpenStudy (anonymous):

they are given to you

OpenStudy (anonymous):

now the third term is tricky, f(n-1) , that is recursive, so it will create several terms in the final sum

OpenStudy (anonymous):

are you with me so far, we have 2 + 4 + f(n-1)

OpenStudy (ashy98):

yes i am (:

OpenStudy (anonymous):

ok so now lets ignore the 2+4 for a minute and focus on f(n-1)

OpenStudy (anonymous):

so what is n?

OpenStudy (ashy98):

2 + 4 + f(n-1) = f(n)=-6?

OpenStudy (anonymous):

no, what is n?

OpenStudy (anonymous):

f(n), we have f(5) so what is n?

OpenStudy (anonymous):

just match the pattern you see given f(n) and f(5)

OpenStudy (anonymous):

the f is the same but the other part is different

OpenStudy (anonymous):

ok another hint, n = 5

OpenStudy (anonymous):

ok so f(5) = f(1) + f(2) + f(n - 1) we already know f(1) and f(2) so we can write as 2 + 4 + f(n - 1) but we also know n so we can write as 2 + 4 + f(5 -1) if we subtract one from the 5 we have f(5) = 2 + 4 + f(4)

OpenStudy (anonymous):

now you need to find f(4), which starts 1 + 2 +...

OpenStudy (ashy98):

Directly substitute n = 5 into the formula. f(5) = f(1) + f(2) + f(4) .. We know f(1) = 2, and f(2) = 4 so f(5) = 2 + 4 + f(4) = 6 + f(4). Now determine f(4) . . . f(4) = f(1) + f(2) + f(3) = 6 + f(3)... so.. f(5) = 6 + (6 + f(3) ) = 12 + f(3) right?

OpenStudy (anonymous):

you got the idea, yep

OpenStudy (anonymous):

so it will be some positive number

OpenStudy (ashy98):

f(5) = 6+16 f(5) = 22

OpenStudy (ashy98):

?

OpenStudy (anonymous):

sounds good to me

OpenStudy (ashy98):

thanks you so much could you help me with one more?

OpenStudy (anonymous):

sure, by the way u see that pattern was 6*(n-2) + f(2) or 6*(n-2) + 4 is the general solution for any n

OpenStudy (ashy98):

yes i did (:

OpenStudy (anonymous):

ok perfect, u understand this very well

OpenStudy (ashy98):

Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2 A.) 2, 3, 5, 7, 9 B.) 2, 3, 4, 7, 11 C.) 2, 3, 5, 9, 11 D.) 2, 3, 5, 8, 13

OpenStudy (anonymous):

ok so yeah slightly more complex but exact same idea, i recommend starting brute force like we did before until you see a general pattern, unless you can just see the pattern right away

OpenStudy (anonymous):

you already know the first two terms which is nice, so you just need f(3), f(4), and f(5)

OpenStudy (anonymous):

so just plug and chug

OpenStudy (ashy98):

C?

OpenStudy (anonymous):

no i dont think so

OpenStudy (anonymous):

so whats f(3)

OpenStudy (anonymous):

aka the third term

OpenStudy (ashy98):

ohhh i new at is D 2, 3, 5, 8, 13

OpenStudy (anonymous):

yeah i beleive thats correct

OpenStudy (ashy98):

thanks again for you time & help!!!(:

OpenStudy (anonymous):

yw :)